the marginal social cost of garbage disposed is: MSCd (gd) = 10, where gd is a 50 gallon drum of garbage disposed. The marginal cost of recycling is: MSCr (gr) = 2gr, where gr is a 50 gallon can of recyclables collected. The marginal benefit for the resident of garbage removed from the residence is: MB(gd, gr) = 800 - 2(gd + gr).

At an economically efficient level, how many [ ] 50-gallon cans of materials a resident shall recycle? (Assuming most of residential waste generated can be recycled, i.e., the resident can recycle as much as desired.)


1
Expert's answer
2021-10-21T09:57:23-0400

Solution:

Marginal social cost of garbage disposed: MSCd (gd) = 10

Marginal cost of recycling: MSCr (gr) = 2gr

Marginal benefit for the resident of garbage removed: MB (gd, gr) = 800 – 2(gd + gr)

At economic efficient level: Marginal benefit = Marginal cost

800 – 2(gd + gr) = 2gr

800 – 2(10 + gr) = 2gr

800 – 20 – 2gr = 2gr

780 = 2gr + 2gr

780 = 4gr

gr = 195

A resident shall recycle 195 - 50gallon cans of materials.